### v2

Abstract
Existing Multimodal Large Language Model (MLLM)-based agents face significant challenges in handling complex GUI (Graphical User Interface) interactions on devices. These challenges arise from the dynamic and structured nature of GUI environments, which integrate text, images, and spatial relationships, as well as the variability in action spaces across different pages and tasks. To address these limitations, we propose MobA, a novel MLLM-based mobile assistant system. MobA introduces an adaptive planning module that incorporates a reflection mechanism for error recovery and dynamically adjusts plans to align with the real environment contexts and action module’s execution capacity. Additionally, a multifaceted memory module provides comprehensive memory support to enhance adaptability and efficiency. We also present MobBench, a dataset designed for complex mobile interactions. Experimental results on MobBench and AndroidArena demonstrate MobA’s ability to handle dynamic GUI environments and perform complex mobile tasks.
[Uncaptioned image] MobA: Multifaceted Memory-Enhanced Adaptive Planning for Efficient Mobile Task Automation
Zichen Zhu, Hao Tang, Yansi Li, Dingye Liu, Hongshen Xu, Kunyao Lan, Danyang Zhang, Yixuan Jiang, Hao Zhou, Chenrun Wang, Situo Zhang, Liangtai Sun, Yixiao Wang, Yuheng Sun, Lu Chen [^1] Corresponding authors., Kai Yu [^1] Corresponding authors. X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, SJTU AI Institute Shanghai Jiao Tong University, Shanghai, China {JamesZhutheThird,chenlusz,kai.yu}@sjtu.edu.cn
1 Introduction
Multimodal large language models (MLLMs) have seen significant advancements in recent years, supported by vast multimodal datasets. These models Hu et al. (2024); Liu et al. (2024a); Ye et al. (2024, 2023); Chen et al. (2024b); Sun et al. (2024a); Liu et al. (2023); Dai et al. (2023); Chen et al. (2023); Zhu et al. (2024); Yao et al. (2024); OpenAI (2023); Team (2024) excel in tasks such as Chain-of-Thought (CoT) reasoning Wei et al. (2022), In-Context Learning (ICL) Brown et al. (2020), and various applications Wang et al. (2024c); Wang and Zhao (2023); Chen et al. (2024a); Liu et al. (2024b); Pan et al. (2024); Ge et al. (2024); Wu et al. (2024); Lee et al. (2024b); Qian et al. (2024b, a). Their capabilities have also enabled new MLLM-based agents for real-world tasks Li et al. (2017, 2019); Sun et al. (2022); Zhu et al. (2023); Zhan and Zhang (2023); Zhang et al. (2023a, 2024a); Nong et al. (2024); Ma et al. (2024); Wang et al. (2024b, a).
{CJK}
UTF8gbsn
However, MLLMs face significant challenges when addressing complex GUI interactions and facing diverse user demands in real-world scenarios, particularly on devices such as smartphones Zhang et al. (2024b) and computers Cao et al. (2024); Xie et al. (2024). On the one hand, GUI environments are highly diverse and pose different action spaces across different apps and pages. For instance, the number and position of clickable icons can vary greatly across pages; some pages require text input, while others involve scrollable elements. Such variability makes proactive task planning hardly adapt to the real environment contexts and thus become infeasible to complete. On the other hand, the action executor can also lack capabilities enough to achieve it, even given a feasible task plan. In all these cases, agents with trivial or static planning Zheng et al. (2024); Zhang et al. (2024a); Nong et al. (2024); Ma et al. (2024); Xing et al. (2024a) will fail to align with the environment contexts and action executor’s capacity and thus can fail the whole task easily caused by failure of a single sub-task. Furthermore, existing MLLM-based GUI agents Zhang et al. (2023b, a); Wang et al. (2024b, a) often lack a powerful and comprehensive memory to face the need for dynamic planning at various levels and diverse user demands. These problems hinder the design of a practical mobile assistant.
To address these challenges, we propose MobA, a novel MLLM-based mobile assistant system with an adaptive planning module that dynamically adjusts task plans according to the execution results. As proactive planning often fails to accurately determine the actions required in a specific application or page or to align with the action executor’s capacity, MobA leverages reflection mechanism to recover task execution from failed sub-plans by reassessing goals or breaking tasks into more fine-grained sub-goals. To better support adaptive planning with various sub-goal granularity, a multifaceted memory module providing hierarchical memory support is proposed. We also introduce MobBench, a diverse dataset for complex mobile interactions, and demonstrate MobA’s effectiveness on MobBench and AndroidArena Xing et al. (2024b), showing its capability to handle dynamic GUI environments.
Our contributions are threefold:
  • We propose an adaptive planning module that incorporates a reflection mechanism for error recovery and dynamically adjusts plans based on the current GUI environment and action executor’s capacity.
  • We develop a multifaceted memory module that provides hierarchical memory support to enhance task adaptability and efficiency.
  • We introduce MobBench, a diverse dataset for complex mobile interactions, and validate the effectiveness of our approach through extensive experiments on MobBench and AndroidArena.
Figure 1: The Illustration of adaptive planning and multifaceted memory structure. There are 4 cases in adaptive planning: (a) Plan reflection failure, indicating the goal needs to be decomposed. (b) In execution reflection failure, the goal needs to be decomposed. (c) Execution reflection failure, the goal needs to be refined. (d) Goal complete.
2 The MobA System
The system overview of MobA is shown in Figure 2.
Figure 2: System Overview of MobA.
{CJK}
UTF8gbsn {CJK}UTF8gbsn {CJK}UTF8gbsn MobA comprises a Global Agent (GA) and a Local Agent (LA). The Global Agent consists of a Plan Module and a Reflection Module. The Plan Module interprets the user’s command (➀) and resolves the task into several easier and clearer sub-tasks adaptively with the help of experiences in the multifaceted Memory Module (➁), while the Reflection Module will review if the decomposition is feasible and sub-goals are achievable. Then, under the direction of a specific sub-goal (➂), the Local Agent will leverage the experiences in the Memory Module (➃), predict the concrete actions, and directly control the device (➄). After LA’s execution, the Reflection Module will reflect if the current sub-task has been completed (➅) and the Plan Module can revise the plan accordingly. The Memory Module can also be updated after the invocation of the Plan Module and Local Agent (➆) to improve MobA’s performance through execution. To ameliorate the performance at the early stage of the memory, it can also be initialized with warm-up of some basic expert experiences. Finally, MobA can generate a response to the user regarding the result of task execution (➇). The remaining parts of this section will elaborate on the proposed adaptive plan module and multifaceted memory module.
2.1 Task Completion with Adaptive Planning
Facing the problem that static fixed-level task planning is deficient in aligning with real environment contexts and the Action Module’s capacity, we propose adaptive planning to react to concrete execution results of the Action Module and adjust the granularity of task decomposition adaptively. The proposed planning workflow is demonstrated in Algorithm 1. Given an established sub-goal, the reflection module is first adopted to review the sub-task feasibility. Then the Action Module will attempt to complete the reviewed sub-goal. The execution result will be inspected again by the reflection module. Once failure is detected, the Plan Module is invoked to revise the task plan to adapt to the current environment context or to further break the sub-goal down to match the Action Module’s execution capacity. By repeating this procedure, MobA can generate a multi-granularity task plan that well aligns with the environment contexts and the Action Module’s capacity iteratively and dynamically.
Input: Global Agent $GA$, Local Agent $LA$, Goal $G_{0}$
task_stack.push($G_{0}$)
while task_stack not empty do
       cur_task $\leftarrow$ task_stack.pop()
      can_do $\leftarrow$ GA.reflect_plan(cur_task)
      if can_do then
             action,$\leftarrow$ LA.exec_task(cur_task)
            cur_task_complete $\leftarrow$ GA.reflect_exec(action,obs)
       end if
      if not can_do or not cur_task_complete then
             new_subtasks $\leftarrow$ GA.plan(cur_task)
            task_stack.push(new_subtasks)
       end if
      GA.updateMemory()
end while
Algorithm 1 Adaptive Planning of MobA
2.2 Multifaceted Memory
The Memory Module serves as the backbone of MobA’s adaptability and learning capabilities, storing historical data to enhance decision-making and reduce redundant actions. It is categorized into five components:
Task Memory: Tracks the execution history of tasks, including task decomposition structures, action traces, success and failure records, and reflections. This hierarchical organization enables efficient retrieval of relevant experiences for task planning and execution.
App Memory: Maintains detailed observations and exploration histories for various applications, including functional descriptions and page-specific interactions. This helps the agent adapt to similar GUI layouts and locate target applications more effectively.
Page Memory: Encompasses the historical steps executed on this interface, such as the positioning of a particular button on the page, among other actions. This facilitates the agent’s ability to perform similar operations on the page based on past interactions more effectively.
Action Memory: Incorporates the operations executed during the current task cycle, enabling the agent to more clearly capture the actions performed within this task and to more precisely define the subsequent steps required.
User Memory: Captures user-specific interaction histories, such as preferences, habitual commands, and implicit requirements. This allows MobA to better infer user intent and personalize task execution.
3 Experiments
To comprehensively compare MobA with other GUI agents in handling complex user instructions and executing GUI interactions on mobile devices, we evaluate them using a real-life scenario test set called MobBench. Additionally, we assess our method using the widely adopted mobile benchmark, Android Arena.
3.1 The MobBench Test Set
The MobBench comprises a diverse test set of 50 tasks designed to evaluate the performance of MobA in real-world mobile application scenarios. The test set includes 10 applications widely used in China, each with four tasks of varying difficulty: Easy, Medium, Hard, and Indirect Comprehension, totaling 40 tasks. The tasks are categorized by the complexity and steps required to complete them. Indirect Comprehension is designed for common cases where the user gives a vague instruction without detailing which application or specific steps are required. The agent is expected to decide target application and find an effective approach. Additionally, there are 10 Cross-Application tasks, which involve interacting with two applications and are more close to Hard level in difficulty. These tasks focus on evaluating the ability of information extraction and retrieval, as well as the awareness of sub-goal completion and application switching.
Compared with several similar task sets mentioned in other papers Zhang et al. (2023a); Wang et al. (2024b, a); Zhang et al. (2024a); Lee et al. (2024a), which only get a score when it finishes the task, we assign several milestone scores for sub-tasks in MobBench. This allows for a more precise process assessment, in the cases where the task is partially finished. We also include a detailed preparation instruction for tasks when a more justice and stable start is needed.
To establish a human expert baseline, three human operators independently perform the tasks on three different mobile phones, documenting their execution steps. The average number of steps taken is used as the human expert baseline.
3.2 Metrics
Three metrics are designed to better compare the capability of GUI agents thoroughly.
Milestone Score (MS): Scoring milestones are assigned to several sub-tasks, evenly distributed during the task completion process. Since each task contains 1 to 6 milestones, the agent will get a score as it reaches each milestone. We sum up all milestone scores of 50 tasks as the primary metric.
Complete Rate (CR): If the agent gets all milestone scores in one task, it is considered as task complete. This is the most common and straightforward metric for GUI agent evaluation.
Execution Efficiency (EE): We record the effective number of steps for each task and the corresponding milestone scores, that is, the total number of steps executed at the time of getting the last effective milestone score, and calculate the average number of steps required to obtain each effective milestone score. The lower this number, the more efficient the execution; the higher it is, the more it includes ineffective actions.
The average milestone scores and execution steps for each task type are summarized in Table 1.
Task Type # Tasks Avg. MS Avg. Steps EE
Easy 10 1.0 4.3 4.30
Medium 10 2.2 7.3 3.32
Hard 10 4.1 15.2 3.71
Indirect 10 2.8 9.4 3.36
Cross-App 10 3.1 10.8 3.48
Overall 50 2.7 9.4 3.56
Table 1: Average scores and expert execution steps for different task types of MobBench.
3.3 Setups
To provide a comprehensive evaluation, MobA is compared against several baselines from basic manual operations to several sophisticated agent-based automation.
Human Baseline as mentioned in § 3.1 are considered as the optimal solution for each task.
GPT-4o + Human Baseline utilizes an iterative process where the GPT model OpenAI (2023) provides guidance for manual task execution.
AppAgent Zhang et al. (2023a) uses both view hierarchy and screenshot for planning and choosing target actions. All interactive elements are marked with bounding boxes and a unique index for better grounding performance.
Mobile Agent (v2) Wang et al. (2024b, a) uses only visual information from screenshots as inputs. Target elements are selected with the guidance of OCR and CLIP Radford et al. (2021) models.
MobA is evaluated under several settings by disabling the Memory Module or/and Plan Module to assess its performance and the impact of these two modules. We disable the Plan Module by replacing the Global Agent with a plain agent, and no sub-tasks are provided to the Action and Reflection Module. We disable the Memory Module by removing all in-context examples and historical experience information (including observations, thoughts, previous actions, and their execution status), focusing on assessing the core capability in zero-shot task execution.
All experiments are conducted using gpt-4o-2024-05-13 API. The primary evaluation metric is the first attempt complete rate, directly measuring the effectiveness of each system in completing tasks on the first try without retries.
3.4 Results and Analysis
The overall experiment results are as listed in Table 2. And for more detailed results categorized by task type please refer to Figure 5.
Model CR MS EE
Human 50/50 133 3.56
GPT-4o + Human 49/50 130 (97.7%) 3.82 (107.2%)
AppAgent 6/50 35 (28.6%) 4.43 (124.4%)
MobileAgent (v2) 17/50 63 (48.9%) 4.84 (136.0%)
MobA w/o M & P 13/50 52 (39.1%) 4.42 (124.2%)
MobA w/o P 15/50 65 (48.9%) 4.17 (117.1%)
MobA w/o M 22/50 72 (54.1%) 3.81 (106.9%)
MobA 28/50 88 (66.2%) 3.44 (96.7%)
Table 2: Overall Performance on MobBench. M: Memory Module. P: Planning Module.
Table 2 shows the performance of four baselines. Due to the complexity of mobile interfaces and the technical limitations encountered during task execution, the overall task completion rates (Complete Rate, CR) are relatively low for all agents. Consequently, the Milestone Score (MS) serves as a finer metric to more accurately reflect the performance of each agent by considering partial task completion. While there are notable differences in Milestone Scores among the baseline models, the gap in Execution Efficiency (EE) is less significant. This is because most agents can smoothly complete simpler sub-goals, whereas, for more complex sub-goals, the agents either complete them or fail entirely, resulting in closer performance regarding execution efficiency.
3.4.1 Performance Comparison
The performance of MobileAgent is notably higher than that of AppAgent. This improvement is mainly due to the inclusion of both Memory and Reflection modules in MobileAgent, which enhance reasoning capacity and utilize more computational resources, such as tokens. Additionally, MobileAgent keeps a record of all historical actions, allowing it to learn from the entire sequence of operations, whereas AppAgent can only track the most recent action. Furthermore, MobileAgent relies on OCR and CLIP modules for target localization, offering greater flexibility and avoiding the technical limitations that AppAgent faces when dependent on XML files. By adopting a twice-reflection strategy, the ineffective execution steps are slightly reduced, where the sub-tasks that are not able to be completed with a single action are decomposed finer before executed. This gives clearer guidance for the Local Agent to decide the target actions.
3.4.2 Ablation Study
The lower part of Table 2 presents the results of the ablation study, where we experimented with four different configurations by selectively enabling or disabling the Memory and Plan modules. The results indicate that incorporating both Memory and Plan modules significantly enhances the agent’s overall performance.
The Plan module alone shows a much stronger effect than the Memory module alone, validating one of the core contributions of this paper—the effectiveness of task decomposition planning. By decomposing tasks into manageable sub-tasks, MobA can perform global planning, avoid redundant actions, and minimize overlooked details, effectively managing its historical actions (since in a tree-structured task, previously completed sub-tasks are inherently tracked). Unlike MobileAgent, which focuses solely on the next specific action, MobA first determines the next abstract task and then plans the specific execution steps, closely mirroring human reasoning patterns and providing a more structured approach.
When the Memory module is introduced, MobA’s performance further improves, particularly in cross-application tasks (see Figure 5 (b)). This enhancement is due to the Memory module’s ability to retain crucial information over longer periods, such as "the day I am traveling to Shenzhen", allowing it to reference previous screens’ key content. In contrast, without the Memory module, the agent is limited to short-term memory of only the current and the immediately preceding steps, resulting in less effective task execution.
3.5 Results on Android Arena
We also performed evaluations on Android Arena Xing et al. (2024b), comprising 157 single-app tasks and 21 cross-app tasks. As shown in Table 3, MobA achieves success rates (SR) of 0.783 on single-app tasks and 0.714 on cross-app tasks, outperforming GPT-4 by 2.4% and 14.3%, respectively. The notable improvement in cross-app tasks is attributed to MobA’s subtask decomposition capability, which enables better app-switching decisions during tasks requiring more steps. Additionally, MobA’s reflection module encourages exploration, reducing repetitive actions and improving task success rates.
The Android Arena evaluation also highlights limitations in task completion judgment with GPT-4, with 11.8% of tasks being misclassified, compared to the results checked by humans. This is partly due to MobA’s tendency to execute redundant actions after completing tasks, complicating GPT-4’s evaluation process. Despite this, MobA’s performance gains emphasize its strength in handling complex multi-step tasks, especially in scenarios requiring extensive exploration and app-switching, as evidenced by the significant improvements in cross-app success rates.
Model SR(single-app) SR(cross-app)
GPT-3.5 0.449 0.048
GPT-4 0.759 0.571
MobA(ours) 0.783 0.714
Table 3: The performance of LLMs and MobA on the Android Arena dataset.
4 Case Study
Figure 3: The Example Case of MobA. Please note that several unimportant stages during the execution of a sub-task are omitted for clarity. The key features for each part are as follows. Task: MobA supports cross-application tasks and can interpret indirect commands. Sub-task 1: Memories are retrieved to select target applications and updated to track the trace. Sub-task 3: MobA will reflect and try other approaches if the attempt is failed. Sub-task 9 and sub-task 13: Memories are used to choose correct actions.
Figure 3 demonstrates how the adaptive planning and multifaceted memory support task completion in MobA. MobA can accurately interpret user intent from command “Help me check when will I reach the travel destination tomorrow." and give decomposed sub-tasks based on historical commands. For sub-task 1, MobA retrieves relevant details from App and User Memory, extracts key information (train schedule and destination), and stores it in Action Memory. When encountering failures, MobA uses historical experiences to reflect and adapt. During sub-task 3, when MobA initially failed to input the train number using the Box_Input function, it reflects on its previous operations and employs a character-by-character input method, completing the task. The key feature of this page will be saved into Page Memory, thus MobA is unlikely to encounter the same failure. Additionally, memory retrieval is crucial for handling contextual tasks. In sub-tasks 9 and 13, although the user doesn’t explicitly specify the travel date or destination in the task request. MobA can rely on previously stored Action Memory data to provide an accurate response.
5 Conclusion and Future Works
This paper presented MobA, an innovative Mobile phone Assistant system empowered by MLLMs. Utilizing a two-level agent structure, comprising a Global Agent and a Local Agent, MobA effectively understands user commands, plans tasks, and executes actions. The combination of Memory and Plan Modules enhances its ability to learn from previous interactions, improving efficiency and accuracy. Our evaluations demonstrated that MobA surpasses existing mobile assistants in handling complex tasks, leveraging multi-level memory, task decomposition, and action-validation mechanisms. These features enable precise task execution even with intricate or indirect commands. Future work will focus on improving the performance on image-only scenarios where the view hierarchy is unattainable, deploying an end-side model on mobile phones for faster response and secured privacy. We hope MobA illustrates the potential of MLLMs-empowered mobile assistants and provides valuable insights for future works.
Appendix A Several Useful Links
Demos of MobA:
Code of MobA:
Prompts used in MobA:
Complete MobBench:
Appendix B Related Work
B.1 LLM Agents
The development of intelligent agents has been significantly influenced by the advancements in large language models (LLMs) and multimodal large language models (MLLMs). LLM-based agents leverage the autonomy, reactivity, proactiveness, and social ability of these models to perceive external environments and make decisions Xi et al. (2023). Emerging abilities, such as chain-of-thought (CoT) reasoning Wei et al. (2022); Wang et al. (2023b); Zhang et al. (2023d) and in-context learning (ICL) Brown et al. (2020); Min et al. (2022). Recent studies have explored LLM-based approaches for reflection Yao et al. (2023); Madaan et al. (2023); Shinn et al. (2023); Xu et al. (2024b), planning Sun et al. (2024b); Qian et al. (2024c); Huang et al. (2024), and memory mechanisms Zhang et al. (2024d, c); Li et al. (2023); Maharana et al. (2024); Lan et al. (2024); Zhang et al. (2023c).
At the same time, the agents that utilize M/LLMs to interact with the environments are quickly developed. These agents possess significantly enhanced capabilities for environment observation, task decomposition, and action decision-making, which enable M/LLMs to solve complex tasks across social simulations Park et al. (2023); Aher et al. (2023); Jo et al. (2023); Lan et al. (2024), embodied robots Wu et al. (2023), software development Qian et al. (2024b, a) and virtual assistants Wang et al. (2023a).
B.2 GUI Agents
B.2.1 Traditional GUI Agents
Controlling graphical user interface (GUI) screens based on user commands is a complex task that involves both GUI understanding and command interpretation. Early approaches to GUI agents focused on embedding and modular systems. For example, several agents Li et al. (2017, 2019) combined natural language and programming demonstrations, allowing users to define tasks via descriptions and demonstrations. This method relied on text and image matching for script-based control of the interface. Traditional GUI agents were largely limited by their reliance on pre-defined rules and manual programming. These agents were effective within controlled environments but struggled with dynamic, real-world GUI contexts due to their lack of flexibility and adaptability. They required specific scripts or rules for each task, making them less robust when handling the diverse and evolving nature of real-world applications.
B.2.2 Advancements with Multimodal Pretrain Models
The advent of multimodal pretraining models Bai et al. (2021); Li et al. (2021b); Li and Li (2023); He et al. (2021); Li et al. (2021a); Wang et al. (2021); Fu et al. (2024) for GUI understanding marked a significant shift in the development of GUI agents. Pretrained agents Sun et al. (2022); Zhu et al. (2023); Zhan and Zhang (2023); Xu et al. (2024a) integrated multimodal information, such as dialogue history, screenshots, and action history, through pretraining. Unlike earlier methods that relied on rigid scripts, these end-to-end models adopted a more human-like approach to interacting with interfaces, enhancing their efficiency in information retrieval and task execution by mapping visual observations and text commands directly into actions.
B.2.3 MLLM-Empowered GUI Agents
The integration of MLLMs in GUI agents has introduced new opportunities to further enhance their capabilities. With the rise of larger scale models, GUI agents Zhang et al. (2023a, 2024a); Lee et al. (2024a) began to leverage advanced reasoning and decision-making processes. These models utilized structural information provided in the view hierarchy (VH) to annotate and locate UI elements, guiding a sequence of atomic actions to achieve specific goals. VH-only agents Wen et al. (2024) depend on the structural information to reason and make decisions, which greatly lowers the cost of inference making it suitable for deployment on the device. Image-only agents Wang et al. (2024b, a); Gao et al. (2024); Yan et al. (2023), which employs optical character recognition (OCR), CLIP Radford et al. (2021) module, and object detection methods to identify operation targets. This image-only approach is particularly effective when the view hierarchy is inaccessible or noisy, but it may also encounter challenges, e.g. opening a target application by clicking when names are hidden, or logos vary across screens.
Appendix C View hierarchy processing
Input: xml file of the current screen
Output: the annotated screen
// First pass: Filter the small elements and all useless attributes
elements $\leftarrow$ (sort(filter(elements), key=area)
[…]
// Second pass: select elements whose overlapping area with former ones is small
foreach element in elements do
       if element is interactive then
             is_valid $\leftarrow$ True foreach selected_element in selected_elements do
                   if overlapping_area is large then
                         is_valid $\leftarrow$ False
                   end if
             end foreach
            if is_valid is True then
                   selected_elements $\leftarrow$ selected_elements + element
             end if
       end if
end foreach
// Third pass: Add the texts and merge the information of text into interactive elements
foreach element in elements do
       foreach selected_element in selected_elements do
             if element is contained in selected_element then
                   merge(element, selected_element)
             end if
       end foreach
end foreach
// Final pass: Sort the elements from left to right, top to bottom
Sort(elements, key=(y,x))
Plot all the interactive elements with their index
Algorithm 2 The Logic of View-Hierarchy Process Algorithm
[…]
Figure 4: An Example Diagram of View-Hierarchy Processing. From left to right are the original image, unprocessed image and processed image. The underlined parts are the properties that are retained after the merge.
Given that (1) large models still exhibit limitations in processing visual information and (2) certain elements of the mobile phone interface cannot be obtained through visual means alone, the view hierarchy (VH) plays a crucial role in enabling agents to effectively interpret the mobile interface. However, the XML files representing mobile interfaces contain a substantial amount of redundant information. This redundancy increases token counts and complicates the agent’s task of identifying key UI elements.
To address this issue, we developed an algorithm designed to filter UI elements. The algorithm consists of four steps: (1) parsing UI elements from the XML file, (2) filtering user-interactable UI elements based on their attributes, and adding them in ascending order of size, unless they exhibit significant overlap with previously added elements, (3) for UI elements containing text, merging the text content with interactive elements if the text is largely contained within those elements, thus enriching the interactive element with explanatory information, and (4) assigning an index to each UI element according to its central coordinates, from left to right and top to bottom, while plain text elements are assigned an index of -1. This ensures that the index ordering aligns more closely with the user’s natural visual scanning behavior.
In summary, the core of our algorithm is the preservation of key interactive elements and their associated textual information, while minimizing occlusion in the image. For example, in the case of the "plane ticket" element demonstrated in Figure 4, the UI element itself does not contain text, and the text information associated with the plane ticket is non-clickable. By merging the two, the agent can infer that clicking the UI element corresponds to selecting the plane ticket.
However, limitations remain in this approach. There are cases where all elements in the XML file are marked as "clickable=false", despite the presence of interactive elements in practice. Additionally, technical limitations sometimes prevent the XML file from accurately reflecting the current state of the interface.
Appendix D Action Space
We provide all actions supported in MobA in Table 4.
Action Type Usage Description
Click single Click(element_index: int) This function clicks the center of the UI element with the specified element index.
Click by Coordinate single Click_by_Coordinate(x: double, y: double) This function simulates a click at the specified x and y coordinates on the screen.
Double Click single Double_Click(element_index: int) This function double clicks the center of the UI element with the specified element index.
Long Press single Long_Press(element_index: int) This function long-presses the center of the UI element with the specified element index.
Scroll single Scroll(element_index: int, direction: str, distance: str or int) This function swipes from the center of the UI element with the specified element index.
Swipe single Swipe(direction: str, distance: str) This function swipes from the center of the screen.
Type single Type(text: str) This function inputs text on the current input box.
Back single Back() This function presses the back key to return to the previous screen or status.
Box Input combination Box_Input(element_index: int, text: str) This function clicks the input box, inputs given text, and confirms it.
Open App system Open_App(description: Optional[str]) This function locates and opens an app with a short description.
Close App system Close_App(package_name: Optional[str]) This function closes the specified app by its package name.
Error system Failed() This function indicates that the task cannot be completed.
Finish system Finish() This function indicates that the task is completed.
Table 4: Available Actions and Descriptions
Appendix E MobBench
We provide five examples of the tasks included in MobBench as shown in Table 5. You can get the complete collection of 50 tasks in both Chinese and English on Huggingface.
Type Application Task Preparation Scoring Milestones Steps
Easy McDonald’s Switch the language of the McDonald’s app to English. Switch to Chinese. 1. Task completion. 6.7
Medium 12306
(China Railway)
Check the schedule for train G104 from Shanghai to Beijing tomorrow, and find out what time it is expected to arrive in Nanjing. - 1. Enter the timetable screen,
2. Correct train number,
3. Task completion.
11.7
Hard Douban Search for the movie "The Shawshank Redemption" on Douban, mark it as "watched", rate it five stars, and leave a positive review. Remove the previous mark, rating, and review of this movie. 1. Correct movie,
2. Correct mark,
3. Correct rating,
4. Positive review.
9.7
Indirect BiliBili If I’m out of mobile data, what videos can I still watch on the phone? Download several videos in advance. 1. Open BiliBili,
2. Check downloads.
3.3
Cross-APP JD.com, WeChat Share the product link of the most recent JD.com order with a WeChat friend, and write a recommendation message. There is an existing order. 1. Enter the order list,
2. Correct order,
3. Suitable message,
4. Task completion.
10.3
Table 5: Several example tasks in MobBench. The content is translated from Chinese.
Appendix F Detailed Results Comparison
While the performance of all models is relatively similar on simpler tasks, MobA demonstrates superior results in more challenging tasks, outperforming other models except for Human and GPT-4o + Human. This suggests that MobA is more efficient in handling complex cases. Additionally, the incorporation of both the Memory Module and Plan Module enhances performance, highlighting their respective contributions to the system’s overall capability.
Figure 5: Performance on MobBench Categorized by Task Type.
F.1 Human is more adaptive and robust to screen interactions
While the human baseline is considered the optimal solution for each task, the GPT-4o + Human method achieves performance very close to that of human operators on all metrics. In the evaluation of GPT-4o + Human, the agent only provides textual task descriptions and an initial screenshot, and the GPT-4o generates detailed step-by-step instructions, which are then executed manually by a human operator.
The eye-catching performance of GPT-4o + Human can be attributed to several factors: (1) a relatively lenient standard in task execution, allowing human operators to interpret GPT-4o’s general instructions flexibly; (2) human operators automatically completing tasks such as OCR, target detection, and localization, ensuring more precise actions; (3) GPT-4o provides a global plan, avoiding redundant or missed steps; (4) technical issues (e.g., inability to retrieve XML files or missing information in the files) do not affect task completion.